# Study on physical properties and characteristics of an anisotropic compact star model using Karmarkar Condition in $\mathcal{F}(\mathcal{Q})$ gravity

### v1

@@ I Introduction
Various universal events  SupernovaSearchTeam:1998fmf ; Planck:2015mvg ; Planck:2015bpv have demonstrated that improvements to general relativity (GR) are essential at a geometric scale. GR, grounded on Riemannian geometry, illustrates the Levi-Civita connection and defines the Ricci curvature $R$ as the essential characteristic of spacetime. This approach omits torsion and non-metricity from the geometry. An alternate approach to characterise GR and Riemannian geometry is teleparallel gravity  aldrovandi2012teleparallel . This theory suggests that the gravitational force originates from torsion $\mathcal{T}$ rather than curvature $R$ in Riemannian geometry. This idea was formally proposed by Einstein, who had earlier utilised teleparallel geometry in his efforts to formulate a unified field theory  einstein2005riemann . A different technique, referred to as non-metricity or symmetric teleparallel gravity, was developed by Nester  nester1998symmetric . This concept suggests that non-metricity $\mathcal{Q}$, which is not impacted by curvature and torsion, acts as the mediator of gravitational interactions. In contrast to General Relativity, where gravity is conveyed through the affine connection instead of the physical manifold, non-metricity is essential in this alternative approach. The Einstein pseudo-tensor, which transforms into a genuine tensor in the geometric framework, is associated with the non-metricity $\mathcal{Q}$. Jimenez et al. BeltranJimenez:2017tkd expanded symmetric teleparallel gravity to align with general relativity, obtaining $F(\mathcal{Q})$ gravity. Given the growing focus on extended gravity theories in modern cosmology, researchers are exploring alternate geometries to achieve a deeper understanding of late cosmic acceleration  Ferraro:2008ey ; Geng:2011aj ; Cai:2015emx ; Jarv:2015odu . Furthermore, it is important to acknowledge that the metric theories may be expanded beyond Riemannian geometry, as shown in  Conroy:2017yln . Harko et al. Harko:2018gxr presented the modified $F(\mathcal{Q})$ gravity by utilizing the metric-affine formalism and an expansion of symmetric teleparallel gravity. Recent research on advanced theories of gravity emphasises the critical examination of the astrophysical and cosmological characteristics of $F(\mathcal{Q})$ gravity. In this way, Hohmann et al. Hohmann:2018wxu investigated the propagation velocity and possible polarizations of gravitational waves within the framework of Minkowski spacetime in $F(\mathcal{Q})$ gravity. Soudi et al. Soudi:2018dhv defined the strong-field characteristics of this gravitational theory by the examination of gravitational wave polarizations. A multitude of studies has investigated $F(\mathcal{Q})$ gravity across diverse contexts, including the analysis of observational data constraints for various parametrizations of $F(\mathcal{Q})$  Lazkoz:2019sjl ; Ayuso:2020dcu , energy conditions  Mandal:2020lyq , cosmography  Mandal:2020buf , bouncing scenarios  Mandal:2021wer ; bajardi2020bouncing , black holes DAmbrosio:2021zpm , and the development of the growth index in matter perturbations, among other issues Nguyen:2021cnb . Furthermore, comprehensive analyses of the $F(\mathcal{Q})$ theory may be found in contemporary literature, including Esposito:2021ect ; Atayde:2021pgb ; Kimura:2020mpk ; Dimakis:2021gby . It is crucial to highlight that $F(\mathcal{Q})$ gravity precisely accounts for the universe’s fast expansion with statistical accuracy comparable to other established modified gravity theories. Therefore, altered gravity theories at an astrophysical scale must be taken in order to deal with this so-called degeneracy. Promising approaches for this purpose include direct monitoring of the galactic core  Mann:2018xkm and the increasing number of gravitational wave findings  LIGOScientific:2016aoc . For a comprehensive examination of this subject, the subsequent sources  errehymy2022anisotropic ; Maurya:2022cyv ; Tangphati:2021tcy include modern contributions across several settings. Stellar evolution describes the ongoing alterations in a star’s interior composition from its starting point to its final extinction. During this transforming process, a star releases a significant quantity of energy as photons or neutrinos. The fascinating properties of these celestial bodies have inspired many scholars to investigate their significance in astronomy and astrophysics. Gravitational collapse is a crucial topic in relativistic astrophysics, intimately linked to the formation of self-gravitating objects. A stable system initially sustains hydrostatic equilibrium until its gravitational forces continuously decrease its size, leading to gravitational collapse. The collapse yields many configurations, including white dwarfs, neutron stars, black holes and gravastars, called compact stars contingent upon the star’s starting mass. The exact solution of an innovative gravastar model derived from the Mazur:2004fk approach within the context of general relativity, particularly through incorporating the cloud of strings and quintessence in paper Javed:2023nls . The composition and features of compact stars have fascinated scientists, leading to inquiries into their many developmental phases and intrinsic properties within the field of gravitational physics. Multiple endeavours have been undertaken to examine the properties of stellar models influenced by diverse physical forces. In this context, the concept of anisotropy is essential for examining the structural processes of compact structures  Herrera:1997plx . Bower and Liang  bondi1992anisotropic examined regionally anisotropic pressure distributions inside spherically symmetric matter, demonstrating that pressure anisotropy significantly influences the parameters defining the hydrostatic equilibrium of stellar systems. Maurya and his associates Maurya:2017uyk . examined compact spherically symmetric systems characterized by anisotropic pressure profiles, providing substantial insights into the stability analysis of small objects. Karamakar and colleagues  paul2015relativistic investigated the influence of anisotropy in cool compact stars. A large number of scholars have exhibited a keen interest in examining the intrinsic and extrinsic characteristics of celestial bodies using gravitational elements. The Karmarkar condition is a crucial framework that establishes a relationship between temporal and radial metric elements, hence improving the examination of compact objects. The Karmakar condition deals with the embedding of a four-dimensional Riemannian manifold into a higher-dimensional manifold. This approach is utilized to investigate the solutions of static spherical spacetime and to analyze the physically viable characteristics of anisotropic stars  Maurya:2016ecj ; Singh:2016nfw . Bhar et al.  Bhar:2017jbl used this technique to investigate the stability requirements of anisotropic compact stars.
@@
Narawade et al. in Narawade:2024pxb reconstructs a cosmological model utilizing covariant $F(\mathcal{Q})$ gravity within Connection-III and FLRW spacetime. The dynamic behaviour of the reconstructed model is thoroughly analyzed using the H(Z) Hubble parameter and various observational datasets. This research offers compelling insights into cosmic acceleration, suggesting that $F(\mathcal{Q})$ gravity can effectively displace the $\Lambda CDM$ model and provides a convincing alternative explanation for the current accelerating expansion of the Universe without relying on the cosmological constant. Using a dynamical system method, Lohakare et al. Lohakare:2024oeu investigated the cosmic stability of $F(\mathcal{Q,B})$ gravity, where $F(\mathcal{Q})$ stands for the nonmetricity scalar and $F(\mathcal{B})$ for the boundary term. According to this study, the $F(\mathcal{Q,B})$ model is a good substitute for the conventional $\Lambda CDM$ model as it provides a strong framework for understanding the dynamics of cosmic expansion and accurately depicts the observed acceleration of the Universe. In  Alwan:2024lng Alwan et al. investigate neutron stars with covariant $F(\mathcal{Q})$ gravity, assuming that the stars are spherically symmetric and static with ideal fluid matter. Numerical techniques employing quadratic, exponential, and logarithmic $F(\mathcal{Q})$ models are used to determine neutron star characteristics including mass, radius, and compactness. A cosmological model within the context of symmetric teleparallel gravity is presented by Narawade et al.in  Narawade:2023rip . They discover a violation of the strong energy requirement, which may offer important insights into the nature of dark energy. In  Narawade:2023tnn , the accelerating cosmological model’s dynamical aspect was examined. Additionally, the model’s evolutionary behaviour of density parameters for the matter-dominated, radiation-dominated, and dark energy phases is demonstrated within the framework of the $F(\mathcal{Q})$ gravity, a modified symmetric teleparallel gravity. In Narawade:2022cgb , a cosmological model of the Universe is presented within the framework of $F(\mathcal{Q})$ gravity, and it is shown that $F(\mathcal{Q})$ gravity offers an alternative to dark energy for explaining the current cosmic acceleration.
In the present study, the main focus was on $F(\mathcal{Q})$ gravity. The modified gravity provides a straightforward composition of the classical GR, which contributes to improving inertial gravitational interaction, without taking into accounting for the affine space-time structure. The special feature of $F(\mathcal{Q})$ gravity is its capacity to interpret the Universe’s late-time acceleration without demanding alterations to scalar fields. Akarsu:2010zm ; BeltranJimenez:2019tme ; Lazkoz:2019sjl . The viability of $F(\mathcal{Q})$ gravity model has been proved in a wide range of domains, including wormhole geometry Harko:2018gxr ; Xu:2019sbp ; DAmbrosio:2020nev ; Maurya:2022vsn ; Maurya:2023kjx ; Banerjee:2021mqk ; Kiroriwal:2023nul and dynamical analysis Lu:2019hra , which covers the cosmological aspects Mandal:2020buf , and has been presented in relation to $F(\mathcal{Q})$ gravity. The mathematical representation of $F(\mathcal{Q})=\mathcal{Q}+m\mathcal{Q}^{n}$ has been used to investigate holographic dark energy Shekh:2021ule which reveals the behaviour of the deceleration parameter changes from negative to positive, indicating that the universe is transitioning from an early phase of deceleration to the current phase of acceleration. In the framework of extended non-metric theories of gravity, Capozziello et al.  Capozziello:2022tvv explored the possibility of cosmological slow-roll inflation. Specifically, they found that standard conformal structure is recovered in a high-energy regime while it breaks in the weak field limit, as demonstrated by generic functions $F(\mathcal{Q})$, particularly $F(\mathcal{Q})=\alpha Q+\beta Q^{m}$. Regardless of the extensively accepted, the teleparallel version of GR shares a formulation with both STGR and GR  Heisenberg:2022mbo . The primary contrast lies in that gravity is allocated to the non-metricity $(\mathcal{Q})$ , whereas curvature and torsion are regarded as zero in STGR. Compatible with the non-metricity and zero value of the torsion restriction, a coincident gauge is always applicable BeltranJimenez:2017tkd . This method is unique for the reason that it eliminates the affine connections. Nevertheless, the development of metrics will differ in different coordinate systems if we continue with the coincident gauge Lu:2019hra and treat the metric as the only fundamental variable in modified STGR theories, such as the $F(\mathcal{Q})$ theory. While research on extended gravity theories is still in its early stages, the symmetric teleparallel, or $F(\mathcal{Q})$, theory of gravity has been receiving a lot of attention. Due to these novel features of $F(\mathcal{Q})$ gravity, several studies have been conducted so far to examine various aspects of the cosmos, including cosmology, energy requirements, covariant formulation, spherical symmetric structure, and signatures that highlight the significance of gravity. The idea behind working in the $F(\mathcal{Q})$ theory of gravity is that on comparing to the $F(\mathcal{R})$ theory, which involves differential equations of the fourth order, the field equations of $F(\mathcal{Q})$ gravity are of the second order Lin:2021uqa , $F(\mathcal{Q})$ gravity extends the theoretical foundation for modelling compact stars by giving new approaches to examining their structure, stability, and observational signature. It offers a promising path for addressing issues concerning severe astrophysical conditions and challenging the limits of our understanding of gravity which is a significant advantage.
@@ II.1 Field equations of $\mathcal{F}(\mathcal{Q})$ gravity
The Levi-Civita affine connection is crucial in explaining and analysing general relativity. When discussed inside a space-time manifold, it aligns with the metric. Various manifolds and affine connections may be utilized, along with diverse theories of gravity. Harada:2020ikm ; BeltranJimenez:2019tme .The Levi-Civita connection identifies nonmetricity $\mathcal{Q}$ and torsion $\mathcal{T}$ as important geometrical concepts alongside curvature. Alleviating these limitations permits the exploration of non-Riemannian geometric theories characterized by non-vanishing curvature, torsion, and nonmetricity. The $\mathcal{F}(\mathcal{Q})$ gravity model was formulated under these circumstances, yielding a gravitational Lagrangian that incorporates an arbitrary nonmetric function ($\mathcal{Q}$). The extension of $\mathcal{F}(\mathcal{Q})$ gravity is extremely significant in the expansion of the Universe. Jimenez et al BeltranJimenez:2019tme was the first to coin the phrase ”symmetric teleparallel gravity,” commonly known as $\mathcal{F}(\mathcal{Q})$ gravity. Based upon the research of  Zhao:2021zab within the framework of $\mathcal{F}(\mathcal{Q})$ gravity, a generic metric spacetime is analyzed, in which the metric tensor $g_{\gamma\Upsilon}$ and the connection coefficients $\Gamma^{\sigma}_{\gamma\Upsilon}$ are treated as distinct entities.The subsequent equation delineates the Nonmetricity of the aforementioned connection:
@@
These three components, in their generic form, are obtained by disassembling this affine relationship into independent components:
@@
The nonmetricity conjugate is defined as:
@@
The independent traces of the preceding equations are as follows:
@@
The $\mathcal{F}(\mathcal{Q})$ gravity is defined by the following action using Lagrange multiplets.:
@@
In Eq. (9), $g$ represents the determinant metric, $\mathcal{F}(\mathcal{Q})$ denotes the nonmetricity of $\mathcal{Q}$, $\mathcal{L}_{n}$ represents the Lagrange density, and $\sigma_{\psi}^{\Psi\gamma\Upsilon}$ represents the Lagrange multipliers. The action given in Eq. (9) has been used to derive the field equations concerning the metric given below
@@
In equation Eq. (10), the subscript $\mathcal{Q}$ denote the derivative of the function $\mathcal{F}(\mathcal{Q})$ with respect to $\mathcal{Q}$, the energy momentum tensor is defined as :
@@
We get the following equations when we change equation Eq. (9) concerning affine connection:
@@
The hypermomentum tensor density is represented by the following equation:
@@
The affine connection without torsion or curvature has the following form:
@@
If we consider a specific coordinate selection that is well-known as a coincident gauge, for which $\Gamma^{\psi}_{\gamma\Upsilon}=0$. After that, the nonmetricity becomes:
@@
The metric for a general static and spherically symmetric space is expressed as :
@@
In the above equation $dw^{2}\equiv d\theta^{2}+sin^{2}{\theta}d{\phi}^{2}$. we get Eq. (19) by putting Eq. (18) in Eq. (8), then $\mathcal{Q}$ is expressed in terms of r,
@@
where prime($\prime$) signifies the derivative with respect to r. An anisotropic fluid with spherically symmetric geometry has the following energy-momentum tensor:
@@
where, $u_{\gamma}$ denotes four-velocity, and $v_{\gamma}$ denotes the unitary space-like vector in the radial direction that fulfils $u^{\gamma}u_{\gamma}=-1,v^{\gamma}v_{\gamma}=1,u^{\gamma}u_{\gamma}=-1,u^{\gamma}v_{\Upsilon}=0$. $\rho(r)$ represents energy density, $p_{r}(r)$ represents the radial pressure in the direction of $v_{\gamma}$, while the orthogonal represents the tangential pressure $(v_{\gamma})$ is denoted by $p_{t}(r)$.The following are the independent components of the equations of motion Eq. (10) for anisotropic fluid Eq. (12) are written as follows:
@@
Assuming a zero affine connection in the coordinate system, $\mathcal{F}(\mathcal{Q})$ gravity Eq. (9) can be expressed as follows:
@@
Combining equation (10) and motion equations yields $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$. This leads us to the conclusion that the function $\mathcal{F}(\mathcal{Q})$ must be linear. Choosing nonlinear values for the function $\mathcal{F}(\mathcal{Q})$ might result in conflicting equations of motion and solutions, especially for $\mathcal{F}(\mathcal{Q}=\mathcal{Q}^{2})$. Applying a nonlinear function of $\mathcal{F}(\mathcal{Q})$ in gravity results in different values for the equation of motion. To analyze and resolve the nonlinear function of $\mathcal{F}(\mathcal{Q})$ gravity, we need a more comprehensive spherically symmetric metric for coincident gauges. Please visit  Zhao:2021zab for a more in-depth study of this topic. In this investigation, we used a linear function with $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$ and found that the spherically symmetric coordinate system (18) corresponds to the affine connection $\Gamma^{\gamma}_{\gamma\Upsilon}=0$. To formulate a realistic compact stellar model, we need to discover a useful function of $\mathcal{F}(\mathcal{Q})$. According to  Wang:2021zaz , matching the compact star solution with the Schwarzchild (Anti-) de sitter solution on the exterior boundary requires selecting a function that assumes $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$. Then, to determine the field equations, we utilise the function $\mathcal{F}(\mathcal{Q})$, which is expressed as follows:

### v2

@@ I Introduction
Various universal events  SupernovaSearchTeam:1998fmf ; Planck:2015mvg ; Planck:2015bpv have demonstrated that improvements to general relativity (GR) are essential at a geometric scale. GR, grounded on Riemannian geometry, illustrates the Levi-Civita connection and defines the Ricci curvature $R$ as the essential characteristic of spacetime. This approach omits torsion and non-metricity from the geometry. An alternate approach to characterise GR and Riemannian geometry is teleparallel gravity  aldrovandi2012teleparallel . This theory suggests that the gravitational force originates from torsion $\mathcal{T}$ rather than curvature $R$ in Riemannian geometry. This idea was formally proposed by Einstein, who had earlier utilised teleparallel geometry in his efforts to formulate a unified field theory  einstein2005riemann . A different technique, referred to as non-metricity or symmetric teleparallel gravity, was developed by Nester  nester1998symmetric . This concept suggests that non-metricity $\mathcal{Q}$, which is not impacted by curvature and torsion, acts as the mediator of gravitational interactions. In contrast to General Relativity, where gravity is conveyed through the affine connection instead of the physical manifold, non-metricity is essential in this alternative approach. The Einstein pseudo-tensor, which transforms into a genuine tensor in the geometric framework, is associated with the non-metricity $\mathcal{Q}$. Jimenez et al. BeltranJimenez:2017tkd expanded symmetric teleparallel gravity to align with general relativity, obtaining $F(\mathcal{Q})$ gravity. Given the growing focus on extended gravity theories in modern cosmology, researchers are exploring alternate geometries to achieve a deeper understanding of late cosmic acceleration  Ferraro:2008ey ; Geng:2011aj ; Cai:2015emx ; Jarv:2015odu . Furthermore, it is important to acknowledge that the metric theories may be expanded beyond Riemannian geometry, as shown in  Conroy:2017yln . Harko et al. Harko:2018gxr presented the modified $F(\mathcal{Q})$ gravity by utilizing the metric-affine formalism and an expansion of symmetric teleparallel gravity. Recent research on advanced theories of gravity emphasises the critical examination of the astrophysical and cosmological characteristics of $F(\mathcal{Q})$ gravity. In this way, Hohmann et al. Hohmann:2018wxu investigated the propagation velocity and possible polarizations of gravitational waves within the framework of Minkowski spacetime in $F(\mathcal{Q})$ gravity. Soudi et al. Soudi:2018dhv defined the strong-field characteristics of this gravitational theory by the examination of gravitational wave polarizations. A multitude of studies has investigated $F(\mathcal{Q})$ gravity across diverse contexts, including the analysis of observational data constraints for various parametrizations of $F(\mathcal{Q})$  Lazkoz:2019sjl ; Ayuso:2020dcu , energy conditions  Mandal:2020lyq , cosmography  Mandal:2020buf , bouncing scenarios  Mandal:2021wer ; bajardi2020bouncing , black holes DAmbrosio:2021zpm , and the development of the growth index in matter perturbations, among other issues Nguyen:2021cnb . Furthermore, comprehensive analyses of the $F(\mathcal{Q})$ theory may be found in contemporary literature, including Esposito:2021ect ; Atayde:2021pgb ; Kimura:2020mpk ; Dimakis:2021gby . It is crucial to highlight that $F(\mathcal{Q})$ gravity precisely accounts for the universe’s fast expansion with statistical accuracy comparable to other established modified gravity theories. Therefore, altered gravity theories at an astrophysical scale must be taken in order to deal with this so-called degeneracy. Promising approaches for this purpose include direct monitoring of the galactic core  Mann:2018xkm and the increasing number of gravitational wave findings  LIGOScientific:2016aoc . For a comprehensive examination of this subject, the subsequent sources  errehymy2022anisotropic ; Maurya:2022cyv ; Tangphati:2021tcy include modern contributions across several settings. Stellar evolution describes the ongoing alterations in a star’s interior composition from its starting point to its final extinction. During this transforming process, a star releases a significant quantity of energy as photons or neutrinos. The fascinating properties of these celestial bodies have inspired many scholars to investigate their significance in astronomy and astrophysics. Gravitational collapse is a crucial topic in relativistic astrophysics, intimately linked to the formation of self-gravitating objects. A stable system initially sustains hydrostatic equilibrium until its gravitational forces continuously decrease its size, leading to gravitational collapse. The collapse yields many configurations, including white dwarfs, neutron stars, black holes and gravastars, called compact stars contingent upon the star’s starting mass. The precise solution of a novel gravastar model, based on the method in Mazur:2004fk , is formulated inside the framework of general relativity, specifically by integrating the cloud of strings and quintessence as presented in Javed:2023nls . The composition and features of compact stars have fascinated scientists, leading to inquiries into their many developmental phases and intrinsic properties within the field of gravitational physics. Multiple endeavours have been undertaken to examine the properties of stellar models influenced by diverse physical forces. In this context, the concept of anisotropy is essential for examining the structural processes of compact structures  Herrera:1997plx . Bower and Liang  bondi1992anisotropic examined regionally anisotropic pressure distributions inside spherically symmetric matter, demonstrating that pressure anisotropy significantly influences the parameters defining the hydrostatic equilibrium of stellar systems. Maurya and his associates Maurya:2017uyk . examined compact spherically symmetric systems characterized by anisotropic pressure profiles, providing substantial insights into the stability analysis of small objects. Karamakar and colleagues  paul2015relativistic investigated the influence of anisotropy in cool compact stars. A large number of scholars have exhibited a keen interest in examining the intrinsic and extrinsic characteristics of celestial bodies using gravitational elements. The Karmarkar condition is a crucial framework that establishes a relationship between temporal and radial metric elements, hence improving the examination of compact objects. The Karmakar condition deals with the embedding of a four-dimensional Riemannian manifold into a higher-dimensional manifold. This approach is utilized to investigate the solutions of static spherical spacetime and to analyze the physically viable characteristics of anisotropic stars  Maurya:2016ecj ; Singh:2016nfw . Bhar et al.  Bhar:2017jbl used this technique to investigate the stability requirements of anisotropic compact stars.
@@
Covariant $F(\mathcal{Q})$ gravity is used by Narawade et al. in Narawade:2024pxb to recreate a cosmological model in Connection-III and FLRW spacetime. Several observational datasets and the H(Z) Hubble parameter are used to extensively evaluate the dynamic behavior of the reconstructed model. This study presents strong evidence for cosmic acceleration, indicating that $F(\mathcal{Q})$ gravity can successfully replace the $\Lambda CDM$ model and makes a strong case for an alternate, non-cosmological constant justification of the Universe’s present accelerating expansion. Using a dynamical system method, Lohakare et al. Lohakare:2024oeu investigated the cosmic stability of $F(\mathcal{Q,B})$ gravity, where $F(\mathcal{Q})$ stands for the nonmetricity scalar and $F(\mathcal{B})$ for the boundary term. According to this study, the $F(\mathcal{Q,B})$ model is a good substitute for the conventional $\Lambda CDM$ model as it provides a strong framework for understanding the dynamics of cosmic expansion and accurately depicts the observed acceleration of the Universe. In  Alwan:2024lng Alwan et al. investigate neutron stars with covariant $F(\mathcal{Q})$ gravity, assuming that the stars are spherically symmetric and static with ideal fluid matter. Numerical techniques employing quadratic, exponential, and logarithmic $F(\mathcal{Q})$ models are used to determine neutron star characteristics including mass, radius, and compactness. A cosmological model within the context of symmetric teleparallel gravity is presented by Narawade et al.in  Narawade:2023rip . They discover a violation of the strong energy requirement, which may offer important insights into the nature of dark energy. In  Narawade:2023tnn , the accelerating cosmological model’s dynamical aspect was examined. Additionally, the model’s evolutionary behaviour of density parameters for the matter-dominated, radiation-dominated, and dark energy phases is demonstrated within the framework of the $F(\mathcal{Q})$ gravity, a modified symmetric teleparallel gravity. In Narawade:2022cgb , a cosmological model of the Universe is presented within the framework of $F(\mathcal{Q})$ gravity, and it is shown that $F(\mathcal{Q})$ gravity offers an alternative to dark energy for explaining the current cosmic acceleration.
This work mostly focused on $F(\mathcal{Q})$ gravity. The modified gravity offers a clear formulation of classical GR, enhancing inertial gravitational interaction while disregarding the affine spacetime framework. The distinctive characteristic of $F(\mathcal{Q})$ gravity is its ability to elucidate the Universe’s late-time acceleration without necessitating modifications to scalar fields Akarsu:2010zm ; BeltranJimenez:2019tme ; Lazkoz:2019sjl . The efficacy of the $F(\mathcal{Q})$ gravity configuration was demonstrated across various domains, involving wormhole geometry Harko:2018gxr ; Xu:2019sbp ; DAmbrosio:2020nev ; Maurya:2022vsn ; Maurya:2023kjx ; Banerjee:2021mqk ; Kiroriwal:2023nul and dynamical analysis Lu:2019hra , encompassing cosmological aspects Mandal:2020buf , and has been articulated in relation to $F(\mathcal{Q})$ gravity. The mathematical expression $F(\mathcal{Q})=\mathcal{Q}+m\mathcal{Q}^{n}$ has been employed to examine holographic dark energy Shekh:2021ule , demonstrating that the deceleration parameter transitions from negative to positive, signifying a shift in the universe via an initial phase of deceleration to the present phase of acceleration. Within the context of extended non-metric theories of gravity, Capozziello et al.  Capozziello:2022tvv investigated the potential for cosmic slow-roll inflation.
They especially discovered that the usual conformal structure is reinstated in a high-energy domain, but it is disrupted in the weak field limit, as shown by generic functions $F(\mathcal{Q})$, notably $F(\mathcal{Q})=\alpha Q+\beta Q^{m}$. Despite widespread acceptance, the teleparallel formulation of General Relativity parallels both STGR and GR  Heisenberg:2022mbo . The main distinction is that gravity is assigned to non-metricity $(\mathcal{Q})$, whereas curvature and torsion are considered negligible in STGR. One may always use a coincident gauge that is consistent for the non-metricity and zero value of the torsion limitation BeltranJimenez:2017tkd .
This approach is distinctive since it removes the affine connections. Nonetheless, the formulation of metrics will vary across distinct coordinate systems if we persist with the coincident gauge Lu:2019hra and regard the metric as the sole essential component in $F(\mathcal{Q})$ theory. Although investigations into extending gravity theories remain still developing, the symmetric teleparallel, or $F(\mathcal{Q})$ gravity has garnered significant interest. Owing to the innovative characteristics of $F(\mathcal{Q})$ gravity, several investigations was undertaken to explore diverse facets of the universe, taking cosmology, energy demands, covariant formulation, spherically symmetric structures, and signatures that underscore the importance of gravity. The premise of the $F(\mathcal{Q})$ theory of gravity is that, in contrast to the $F(\mathcal{R})$ theory, which entails fourth-order differential equations, the field equations of $F(\mathcal{Q})$ gravity are second-order Lin:2021uqa . This theory enhances the theoretical framework for modeling compact stars by providing novel methodologies for analyzing their structure, stability, and observational characteristics. It offers a viable avenue for tackling difficulties related to extreme astrophysical circumstances and testing the boundaries of our comprehension of gravity, which is a notable consequence.
@@ II.1 Field equations of $\mathcal{F}(\mathcal{Q})$ gravity
The Levi-Civita affine connection is crucial in explaining and analysing general relativity. When discussed inside a space-time manifold, it aligns with the metric. Various manifolds and affine connections may be utilized, along with diverse theories of gravity. Harada:2020ikm ; BeltranJimenez:2019tme .The Levi-Civita connection identifies nonmetricity $\mathcal{Q}$ and torsion $\mathcal{T}$ as important geometrical concepts alongside curvature. Alleviating these limitations permits the exploration of non-Riemannian geometric theories characterized by non-zero torsion, nonmetricity, and curvature. The $\mathcal{F}(\mathcal{Q})$ gravity model has been established under these circumstances, resulting in a gravitational Lagrangian that includes an undefined non-metricity ($\mathcal{Q}$). The generalization of $\mathcal{F}(\mathcal{Q})$ gravity is very consequential for the expanding of the Universe. Jimenez et al. BeltranJimenez:2019tme were the pioneers in introducing the term ”symmetric teleparallel gravity,” often referred to as $\mathcal{F}(\mathcal{Q})$ gravity. Based upon the research of  Zhao:2021zab within the framework of $\mathcal{F}(\mathcal{Q})$ gravity, a generic metric spacetime is analyzed, in which the metric tensor $g_{\gamma\Upsilon}$ and the connection coefficients $\Gamma^{\sigma}_{\gamma\Upsilon}$ are treated as distinct entities.The subsequent equation delineates the Nonmetricity of the aforementioned connection:
@@
where the affine connection is:
@@
The nonmetricity conjugate is characterized as follows:
@@
The independent derivations of the aforementioned equations are outlined below:
@@
The $\mathcal{F}(\mathcal{Q})$ gravity is characterized by a specific action via Lagrangian distribution:
@@
In Eq. (9), $g$ signifies the determinant metric, $\mathcal{F}(\mathcal{Q})$ indicates the nonmetricity scalar $\mathcal{Q}$, and $\mathcal{L}_{n}$ specifies the Lagrange density, and $\sigma_{\psi}^{\Psi\gamma\Upsilon}$ represents the Lagrange multipliers. The field equations pertaining to the metric below have been derived using the action provided in Eq. (9) as:
@@
where, $\mathcal{F}_{\mathcal{Q}}=\frac{\partial\mathcal{F}}{\partial{\mathcal{Q}}}$, and The energy-momentum tensor (EMT) can be described as:
@@
We get the below system upon modifying equation Eq. (9) related to affine connection:
@@
The hypermomentum tensor density is given as:
@@
In absence of torsion or curvature, the affine connection takes the form:
@@
When we examine a certain coordinate choice recognized as a coincident gauge, for which $\Gamma^{\psi}_{\gamma\Upsilon}=0$. After that, the nonmetricity becomes:
@@
The metric for a generic static and spherically symmetric (SS) spacetime is articulated as:
@@
In the above equation $dw^{2}\equiv d\theta^{2}+sin^{2}{\theta}d{\phi}^{2}$. Equation  (19) is derived by substituting Equation  (18) into Equation  (8), then $\mathcal{Q}$ is expressed in terms of r,
@@
where the derivative with respect to r is denoted by prime($\prime$) and $\mathcal{T}_{\gamma\Upsilon}$ is given as:
@@
where, $u_{\gamma}$ denotes four-velocity, and $v_{\gamma}$ denotes the unitary space-like vector in the radial direction that fulfils $u^{\gamma}u_{\gamma}=-1,v^{\gamma}v_{\gamma}=1,u^{\gamma}u_{\gamma}=-1,u^{\gamma}v_{\Upsilon}=0$. Energy density is represented by $\rho(r)$, radial pressure in the direction of $v_{\gamma}$ is represented by $p_{r}(r)$, and tangential pressure is represented by the orthogonal. $p_{t}(r)$ represents $(v_{\gamma})$.In the equations of motion Eq. (10) for anisotropic fluid Eq. (12), each of the independent variables are expressed as follows:
@@
With the assumption of a zero affine connection in the coordinate system, the gravity equation $\mathcal{F}(\mathcal{Q})$ can be stated as follows:
@@
Combining equation (10) and motion equations yields $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$. This implies that the function $\mathcal{F}(\mathcal{Q})$ have to be linear. Selecting nonlinear form of $\mathcal{F}(\mathcal{Q})$ may not satisfy field equations and its solutions, particularly when $\mathcal{F}(\mathcal{Q}=\mathcal{Q}^{2})$ is used. The nonlinear function of $\mathcal{F}(\mathcal{Q})$ gravity requires an extensive SS metric for coincident gauges in order to be analyzed and resolved  (see Zhao:2021zab ). In this study, we employed a linear form of $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$ and determined that the SS coordinate system (18) aligns well with $\Gamma^{\gamma}_{\gamma\Upsilon}=0$. In order to develop a compact stellar model that is realistic, it is necessary to identify a appropriate function of $\mathcal{F}(\mathcal{Q})$. The choice of a function that asserts $\mathcal{F}_{\mathcal{Q}\mathcal{Q}}=0$ is essential to align the our metric with the exterior Schwarzschild (Anti-) de Sitter metric at the external boundary, according to  Wang:2021zaz . The function $\mathcal{F}(\mathcal{Q})$ is then employed to ascertain the field equations that can be defined as follows: